# Web Agents with World Models: Learning and Leveraging Environment Dynamics in Web Navigation

### v1

@@ Abstract
Large language models (LLMs) have recently gained much attention in building autonomous agents. However, performance of current LLM-based web agents in long-horizon tasks is far from optimal, often yielding errors such as repeatedly buying a non-refundable flight ticket. By contrast, humans can avoid such an irreversible mistake, as we have an awareness of the potential outcomes (e.g., losing money) of our actions, also known as the “world model”. Motivated by this, our study first starts with preliminary analyses, confirming the absence of world models in current LLMs (e.g., GPT-4o, Claude-3.5-Sonnet, etc.). Then, we present a World-model-augmented (WMA) web agent, which simulates the outcomes of its actions for better decision-making. To overcome the challenges in training LLMs as world models predicting next observations, such as repeated elements across observations and long HTML inputs, we propose a transition-focused observation abstraction, where the prediction objectives are free-form natural language descriptions exclusively highlighting important state differences between time steps. Experiments on WebArena and Mind2Web show that our world models improve agents’ policy selection without training and demonstrate superior cost- and time-efficiency compared to recent tree-search-based agents. The codes are publicly available at https://github.com/kyle8581/WMA-Agents.
@@ 4.2 Inference-time Policy Optimization with the World Model
We begin by sampling $k$ action candidates $\{a_{t}^{1},a_{t}^{2},...,a_{t}^{k}\}$ from $\theta$ via top-$p$ decoding (Holtzman et al., 2019), explore diverse next states $s_{t+1}\in\mathcal{S}$ (Wang et al., 2022). Next, we use the world model $\phi$ to “simulate” the potential next observation $\tilde{o}_{t+1}^{i}$ caused by each action candidate $a_{t}$:
@@ 5.1 Setups and Implementation Details
No Policy Opt. + Policy Opt.
@@ 6.1 Combining Self-refine with Our World Models
Besides our inference-time policy optimization, another way of using world models to improve the policy model is prompting it to refine its predicted action (Madaan et al., 2024), based on the outcome simulated by the world model. Such self-refinement has been showing promising performance in diverse LLM applications (Shinn et al., 2024; Chae et al., 2024).
Here, we conduct a demonstrative experiment of combining self-refine with our world model in the Map domain from WebArena. Since tasks in this domain involve a complex set of utility tools, such as sorting and zoom-in, we consider it suitable for testing self-refine. In this experiment, after the policy model $\theta$ produces an action $a_{t}$ based on the current observation $o_{t}$, we use our world model to simulate the next observation $\tilde{o}_{t}$ and prompt $\theta$ to refine the action based on $\tilde{o}_{t}$. Simply put, this setting allows $\theta$ to make adjustments to its output action when the predicted next observation is not optimal. Table 7 shows that refining with simulated environment feedback improves the agent’s policy selection by 1.8% percentage point in terms of accuracy compared to CoT. While this provides a plausible direction for future work, our simulate-score-select paradigm (Figure 3 (bottom)) yields a almost 2x higher accuracy (13.4% vs. 22.3%), making it our choice of the policy optimization method.666 As a pilot/demonstrative trial, we do not consider sampling multiple actions for self-refinement.
@@ 7 Conclusions
Appendix A Limitations
This paper has the following limitations:
Single modality.
This study focuses on building text-based world models and web agents. In web navigation, however, visual information can also play a critical role (Liu et al., 2024; Zheng et al., 2024). Although HTML and accessibility tree do represent the visual structure of a webpage to some degree, it is recommended for future work to incorporate visual information in addition to textual information for improving the learning of dynamics in the environment (Koh et al., 2024).
@@ Appendix A Limitations
Our current approach demonstrates significant performance gains in selecting actions within a single time step, i.e., from time $t$ to time $t+1$. However, we believe that our idea can be extended to multi-step planning (i.e., predicting $o_{t+n}$ where $n>1$). This can be achieved by recursively feeding the predicted next state back into the world model as the new observation, along with action candidates generated at each time step. Incorporation with techniques such as Monte Carlo Tree Search (Koh et al., 2024) or other planning algorithms is also feasible. Future work can explore such multi-step planning to enable web agents to understand longer-term consequences of their actions and make more informed decisions in scenarios where planning multi-step ahead is necessary.

### v2

@@ Abstract
Large language models (LLMs) have recently gained much attention in building autonomous agents. However, performance of current LLM-based web agents in long-horizon tasks is far from optimal, often yielding errors such as repeatedly buying a non-refundable flight ticket. By contrast, humans can avoid such an irreversible mistake, as we have an awareness of the potential outcomes (e.g., losing money) of our actions, also known as the “world model”. Motivated by this, our study first starts with preliminary analyses, confirming the absence of world models in current LLMs (e.g., GPT-4o, Claude-3.5-Sonnet, etc.). Then, we present a World-model-augmented (WMA) web agent, which simulates the outcomes of its actions for better decision-making. To overcome the challenges in training LLMs as world models predicting next observations, such as repeated elements across observations and long HTML inputs, we propose a transition-focused observation abstraction, where the prediction objectives are free-form natural language descriptions exclusively highlighting important state differences between time steps. Experiments on WebArena and Mind2Web show that our world models improve agents’ policy selection without training and demonstrate superior cost- and time-efficiency compared to recent tree-search-based agents.
@@ 4.2 Inference-time Policy Optimization with the World Model
We begin by sampling $k$ action candidates $\{a_{t}^{1},a_{t}^{2},...,a_{t}^{k}\}$ from $\theta$ via top-$p$ decoding (Holtzman et al., 2019), to conduct diverse exploration on future observations $\{o_{t+1}^{1},o_{t+1}^{2},...,o_{t+1}^{k}\}$ similar to Koh et al. (2024). Next, we use the world model $\phi$ to “simulate” the potential next observation $\tilde{o}_{t+1}^{i}$ caused by each action candidate $a_{t}$:
@@ 5.1 Setups and Implementation Details
w/o Action Selection w/ Action Selection
@@ 6.1 Combining Self-refine with Our World Models
Besides our inference-time policy optimization, another way of using world models is to refine its predicted action (Madaan et al., 2024), based on the outcome simulated by the world model. Such self-refinement has been showing promising performance in diverse LLM applications (Shinn et al., 2024; Chae et al., 2024). Here, we conduct a demonstrative experiment of combining self-refine with our world model in the Map domain from WebArena. Since tasks in this domain involve a complex set of utility tools, such as sorting and zoom-in, we consider it suitable for testing self-refine. In this experiment, after the policy model $\theta$ produces an action $a_{t}$, we use our world model to simulate the next observation $\tilde{o}_{t+1}$ and prompt $\theta$ to refine the action based on $\tilde{o}_{t}$. Simply put, this setting allows $\theta$ to make adjustments to its output action when the predicted next observation is not optimal. Table 7 shows that refining with simulated environment feedback improves the agent’s policy by 1.8% point in terms of accuracy compared to CoT. While this is a plausible direction for future work, our simulate-score-select paradigm yields an almost 2x higher accuracy, making it our choice of the policy optimization method.
@@ 7 Conclusions
Acknowledgments
This work was supported by STEAM R&D Project, NRF, Korea (RS-2024-00454458) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00457882, National AI Research Lab Project). Jinyoung Yeo is the corresponding author.
@@ Acknowledgments
Appendix A Further Analyses
Extending our world model to take multimodal input.
This study focuses on building text-based world models and web agents. In web navigation, however, visual information can also play a critical role (Liu et al., 2024; Zheng et al., 2024). Although HTML and accessibility tree do represent the visual structure of a webpage to some degree, it is helpful to use visual information in addition to textual information for improving the learning of dynamics in the environment (Koh et al., 2024). Thus, we extend our world model to a multimodal setting, inspired by the recent success of multimodal web agents. For our experiments, we use the Mind2Web (cross-task) dataset with Qwen2-VL-2B as the backbone Vision Language Model.
The results are shown in Table 8. Despite using a smaller parameter size compared to Llama-3.1-8B, the multimodal input leads to notable improvements across all metrics. These results demonstrate two key findings: (1) our framework can readily adapt to multimodal settings, and (2) the addition of visual modality provides clear benefits. Given that we used a naive approach to image input, we expect further improvement when incorporating more sophisticated image prompting techniques, such as SeeAct or Set-of-Marks, which could further enhance performance.
Table 8: Comparison of multimodal and text-only world models on Mind2Web (cross-task).
Method Modality of WM EA AF1 Step SR SR
MindAct - - - 17.4 0.8
AWM - 78.3 74.1 62.8 15.3
AWM+WMA (Llama-3.1-8B) Text 79.9 75.8 67.0 25.4
AWM+WMA (Qwen2-VL-2B) Text 79.2 75.1 65.0 23.7
AWM+WMA (Qwen2-VL-2B) Text+Image 83.0 78.9 72.8 36.7
@@ Appendix A Further Analyses
Beyond searching for the best action in a single step, we also explore finding optimal paths through multi-step search by iteratively using the world model. However, prediction errors accumulate as search depth increases when using only world model simulations, limiting real-world applicability. To address this, we adopt a hybrid approach combining simulated interactions for width exploration and actual environmental interactions for depth exploration. We use an A*-like search algorithm following Koh et al. (2024) and conduct experiments using the same settings as shown in Figure 6 (Ablation on the number of sampled actions).
Table 9 shows the results. We find that increasing the depth from $(w=3,d=1)$ to $(w=3,d=3)$ improves performance (17.1 $\rightarrow$ 19.6). However, when comparing settings with the same exploration budget—$(w=9,d=1)$ vs. $(w=3,d=3)$—allocating more budget to width shows slightly better performance. A specific challenge explains this issue: Many errors occur during the execution of optimal action sequences due to mismatches between search time and execution time. Even when a browser loads the same content from the same URL, element IDs calculated by the backend can change upon page reload. This means that the same element might have different IDs between the search and execution phases.
Table 9: Success rates for different width ($w$) and depth ($d$).
Width ($w$) / Depth ($d$) $\mathbf{d=1}$ $\mathbf{d=2}$ $\mathbf{d=3}$
$w=1$ 7.1 - 10.7
$w=2$ 10.1 12.6 -
$w=3$ 17.1 - 19.6
$w=9$ 20.5 - -
In-depth analysis on the world model.
We conduct in-depth analyses to evaluate how well our world model predicts the next observation, focusing primarily on coverage. Coverage measures how much information from the ground-truth observation is successfully captured in the model’s prediction. We define coverage as the ratio of the information of the ground-truth observation covered by the predicted next observation. Specifically, the coverage score is calculated as:
Coverage=# sentences in ground-truth observation covered by the predicted observation# total sentences in ground-truth $\text{Coverage}=\frac{\text{\# sentences in ground-truth observation covered by the predicted observation}}{\text{\# total sentences in ground-truth observation}}.$ (4)
For evaluation, we employ an LLM-as-a-judge approach using GPT-4o as the judge LLM. We begin by separating both predicted and ground-truth observations into sentences. Then, we use an LLM to determine whether information from the target sentence can be found in the source text, which consists of a list of sentences. We run the evaluation on 100 samples used in our preliminary analysis and compare our world model with a few different LLMs, including GPT-4o-mini and GPT-4o.
Table 10: Coverage comparison of different approaches of implementing world models.
Model Coverage (%)
GPT-4o-mini 33.50
GPT-4o 33.85
Ours 42.99